The length of the hypotenuse of 30-60-90 a triangle is 4. What is the perimeter?

Answer by stanbon(75887)

The length of the hypotenuse of 30-60-90 a triangle is 4. What is the perimeter?
----------------------------
If side opposite 30 deg = 1
Then hypotenuse = 2
And 3rd side = sqrt(2^2-1^2) = sqrt(3)
-------------------------------
Your Problem::
If hypot = 4, then side opposite 30 = 2, and 3rd side is 2sqrt(3)
----
Perimeter = 4 + 2 + 2sqrt(3) = 6 + 2sqrt(3)
----------------
Cheers,
Stan H.
------------